# Near-Field Integrated Sensing and Communications for Secure UAV Networks

### v1

@@ Abstract
Integrated sensing and communications (ISAC), secure communications, near field, unmanned aerial vehicle (UAV80), 3D velocities sensing
@@ I Introduction
With the fast development of the low-altitude economy, unmanned aerial vehicles (UAVs) have received significant attention for potential usage in military, civilian, and commercial applications [1]. Particularly, benefiting from the high mobility and the low operation cost, UAVs are anticipated to play significant roles in future wireless networks [2]. In contrast to the terrestrial communications, the ground-to-air communications have higher probability to be dominated by the line-of-sight (LoS) links [3], which is beneficial for improving the communication quality. Moreover, the UAV’s locations can be flexibly adjusted to adapt to the time-varying communications environment, which helps to construct preferable channel conditions in real time.
Despite the above advantages, one prominent challenge for facilitating UAV communications is the potential eavesdropping risk posed by the inherent broadcast nature of the LoS-dominated channels [4]. As such, it is critical to develop physical layer security (PLS) technologies to provide flexible security services for safeguarding UAV communications. Thanks to the capability of the multiple-input multiple-output (MIMO) technique for steering the signals toward the desired directions, the secrecy rate at the receivers can be enhanced via effective beamforming [5, 6]. However, since the eavesdroppers are usually non-cooperative nodes, it is impractical to obtain the wiretap channel state information (CSI) with the conventional pilot-based estimation approach [7]. In addition, due to the high mobility of the UAV, the CSI tracking is a tough task to guarantee stable communication performance while meeting the strict latency requirement. Recently, the concept of sensing-aided communication, which leverages the integrated sensing and communications (ISAC) techniques [8], has emerged for obtaining the communication/wiretap CSI and thereby assisting the subsequent beamforming design [9, 10]. This approach eliminates the requirement for transmitting pilot signals, and instead estimating the CSI based on the distance and direction parameters obtained from echo signals reflected by legitimate users/eavesdroppers [11].
@@ I-B Motivations and Contributions
  • For communications, near-field propagation enables the unique beamfocusing capability, i.e., targeting the signal on a specific location in the near field instead of a specific direction in the far field. This provides enhanced flexibility to prevent the legitimate information leakage and thus anti-eavesdropping. This is beneficial for secure UAV communications.
@@ I-C Organization and Notations
[…]
@@ II System Model and Problem Formulation
As shown in Fig. 1, we consider a near-field UAV communication system that consists of one ISAC GBS, one C-UAV receiving signals from the GBS, and one E-UAV wiretapping the GBS information. The GBS is deployed with a rectangular uniform planar array (UPA) that consists of $M=M_{x}\times M_{y}$ antennas with the spacing $d$, while the C-UAV and the E-UAV are both equipped with a single antenna. The index of the $m$-th GBS antenna is given by $m=aM_{y}+b$, where $a=0,\dots,M_{x}-1$ and $b=1,\dots,M_{y}$. We assume that the GBS operates in the full-duplex mode, which allows it to transmit and receive signals simultaneously. With the aid of advanced self-interference mitigation techniques [31], we assume that the full-duplex GBS can realize perfect self-interference mitigation. The bandwidth of the system, denoted by $B$, is related to the symbol duration $T_{s}=\frac{1}{B}$. Let $N$ denote the number of symbol durations in a single coherent processing interval (CPI) with the duration of $\Delta t$ for the near-field sensing, during which the UAVs’ location and velocity are assumed to be unchanged. We set a Cartesian coordinate system for which the origin is chosen as the center of the GBS UPA and the $x$ and $y$ axes are aligned with the column and row of the UPA, respectively, while the $z$ axis points up to complete a right handed coordinate frame. As such, the coordinate of the UPA center is given by $\mathbf{q}_{b}=[0,0,0]^{T}$, while that of the $m\mathrm{\text{-}th}$ antenna is denoted by $\mathbf{q}_{m}=[x_{m},y_{m},0]^{T}$. In the $l\in\{1,...,L\}\mathrm{\text{-}th}$ CPI, the position and the velocity of the E-UAV are denoted by $\mathbf{q}_{e,l}=\left[x^{e}_{l},y^{e}_{l},z^{e}_{l}\right]^{T}$ and $\mathbf{v}_{e,l}=[{v}^{x}_{e,l},{v}^{y}_{e,l},{v}^{z}_{e,l}]^{T}$, respectively, and the position of the C-UAV is represented by $\mathbf{q}_{c,l}=\left[x^{c}_{l},y^{c}_{l},z^{c}_{l}\right]^{T}$. Given that the E-UAV trajectory is unknown at the GBS, the GBS sends artificial noise (AN) signals to the E-UAV for both sensing and jamming purpose. Then, the transmitted signal $\mathbf{x}_{l}(n)$ of the GBS in the time slot $n$ of the $l$-th CPI is given by
@@ II-B Near-Field Communication Signal Model
where $r^{c}_{l},\theta^{c}_{l},\phi^{c}_{l}$ represent the distance, azimuthal angle, and elevation angle between the C-UAV and the GBS UPA center in the $l\text{-}\mathrm{th}$ CPI, respectively. Accordingly, the achievable data rate at the C-UAV in the $l$-th CPI is given by
@@
$\mathbf{h}^{be}_{l}=\frac{\beta_{0}}{r_{r}^{e}}\boldsymbol{\alpha}(r^{e}_{l},\theta^{e}_{l},\phi^{e}_{l}).$ (11)
@@ III E-UAV 3D Trajectory Tracking Scheme
where $\mathbf{X}_{l}\left(\mathbf{f}^{D}_{l}\right)=\left[\mathbf{G}_{l}(1)\mathbf{x}_{l}(1),\dots,\mathbf{G}_{l}(N)\mathbf{x}_{l}(N)\right]$, $\mathbf{G}_{l}(n)=\mathbf{A}(r^{e}_{l},\theta^{e}_{l},\phi^{e}_{l})\odot\mathbf{D}_{l,n}(\mathbf{f}_{l}^{D})$, and $\mathbf{Z}_{l}^{b}=\left[\mathbf{z}_{l}(1),\mathbf{z}_{l}(2),\ldots,\mathbf{z}_{l}(N)\right]$. As shown in Fig. 2, to facilitate the 3D velocities derivation, we denote the E-UAV’s velocity in the spherical coordinate system that is centered at the GBS UPA center, as $\tilde{\mathbf{v}}_{e,l}=\left[v_{e,l}^{r},v_{e,l}^{\theta},v_{e,l}^{\phi}\right]$, where $v_{e,l}^{r}$, $v_{e,l}^{\theta}$, and $v_{e,l}^{\phi}$ represent the radial velocity, the azimuthal velocity, and the elevation velocity, respectively. As shown in Fig. 3, the mathematical relationship between $\mathbf{f}_{l}^{D}$ and $\tilde{\mathbf{v}}_{e,l}$ is expressed as
@@
$v^{r}_{e,l,m}$, $v^{\theta}_{e,l,m}$, and $v^{\phi}_{e,l,m}$ are respectively given by
@@
where $\alpha_{r,m},\alpha_{\theta,m}$ and $\alpha_{\phi,m}$ are the angles as labeled in Fig. 3.
@@
With the relationship between $\mathbf{f}_{l}^{D}$ and $\tilde{\mathbf{v}}_{e,l}$, the ML method can be utilized for the estimation of $\tilde{\mathbf{v}}_{e,l}$ [30]. Specifically, the ML optimization problem is formulated as follows:
@@
$\hat{\beta}=\frac{\operatorname{tr}\left(\mathbf{Y}^{b}_{l}\left(\mathbf{X}_{l}^{H}\left(\tilde{\mathbf{v}}_{e,l}\right)\right)\right)}{\|\mathbf{X}_{l}\left(\tilde{\mathbf{v}}_{e,l}\right)\|_{F}^{2}}.$ (18)
@@
$\hat{\mathbf{v}}_{e,l}=\arg\max_{\tilde{\mathbf{v}}_{e,l}}g\left(\tilde{\mathbf{v}}_{e,l}\right)=\frac{\left|\mathrm{tr}\left(\mathbf{Y}^{b}_{l}\left(\mathbf{X}_{l}^{H}\left(\tilde{\mathbf{v}}_{e,l}\right)\right)\right)\right|^{2}}{\|\mathbf{X}_{l}\left(\tilde{\mathbf{v}}_{e,l}\right)\|_{F}^{2}}.$ (19)
@@
×$\times\left[d_{l,n,1}\left(\tilde{\mathbf{v}}_{e,l}\right)\frac{\partial v_{e,l,1}}{\partial v^{i}_{e,l}},\ldots,d_{l,n,M}\left(\tilde{\mathbf{v}}_{e,l}\right)\frac{\partial v_{e,l,M}}{\partial v^{i}_{e,l}}\right]^{T},\vphantom{\int}$
@@
where $\tilde{\mathbf{v}}_{e,l}^{(t)}$ and $\eta^{(t)}$ denote the updated value of $\tilde{\mathbf{v}}_{e,l}$ and the step size in the $t$-th iteration. The details of the gradient-descent algorithm is shown in Algorithm 1.
@@ III-B1 State Measurement Error Model
Assuming that the measurement errors for each state are uncorrelated, we can construct the measurement error covariance matrix as $\mathbf{P}_{l}=\text{diag}\{\sigma^{2}_{r^{e}_{l}},\sigma^{2}_{{\theta}^{e}_{l}},\sigma^{2}_{{\phi}^{e}_{l}},\sigma^{2}_{v^{r}_{e,l}},\sigma^{2}_{v^{\theta}_{e,l}},\sigma^{2}_{v^{\phi}_{e,l}}\}$. It has been demonstrated in [34] that the measurement errors are inversely proportional to the signal-to-noise ratio (SNR) of the echo signals. Therefore, we have $\sigma^{2}_{r^{e}_{l}}=\frac{c_{r}}{\mathrm{SNR}_{l}}$, $\sigma^{2}_{\theta^{e}_{l}}=\frac{c_{{\theta}}}{\mathrm{SNR}_{l}}$, $\sigma^{2}_{\phi^{e}_{l}}=\frac{c_{{\phi}}}{\mathrm{SNR}_{l}}$, $\sigma^{2}_{v_{r,l}}=\frac{c_{v_{r}}}{\mathrm{SNR}_{l}}$ , $\sigma^{2}_{v_{\theta}}=\frac{c_{v_{\theta}}}{\mathrm{SNR}_{l}}$ and $\sigma^{2}_{v_{\phi,l}}=\frac{c_{v_{\phi}}}{\mathrm{SNR}_{l}}$, where $c_{r}$, $c_{{\theta}}$, $c_{{\phi}}$, $c_{v_{r}}$,$c_{v_{\theta}}$ and $c_{v_{\phi}}$ are the scaling factors [35], while $\mathrm{SNR}_{l}$ is given by
@@ III-B2 State Prediction Model
where $\boldsymbol{\zeta}_{l}=\left[\zeta_{x^{e}_{l}},\zeta_{y^{e}_{l}},\zeta_{z^{e}_{l}},\zeta_{v^{x}_{e,l}},\zeta_{v^{y}_{e,l}},\zeta_{v^{z}_{e,l}}\right]^{T}$, $\zeta_{x^{e}_{l}},\zeta_{y^{e}_{l}},\zeta_{z^{e}_{l}},\zeta_{v^{x}_{e,l}},\zeta_{v^{y}_{e,l}}$, and $\zeta_{v^{z}_{e,l}}$ denote the prediction errors for $x^{e}_{l},y^{e}_{l},z^{e}_{l},v^{x}_{e,l},v^{y}_{e,l}$, and $v^{z}_{e,l}$ respectively, with $\zeta_{x^{e}_{l}}\sim\mathcal{N}(0,\sigma_{x^{e}_{l}}^{2}),\zeta_{y^{e}_{l}}\sim\mathcal{N}(0,\sigma_{y^{e}_{l}}^{2}),\zeta_{z^{e}_{l}}\sim\mathcal{N}(0,\sigma_{z^{e}_{l}}^{2}),\zeta_{v^{x}_{e,l}}\sim\mathcal{N}(0,\sigma_{v^{x}_{e,l}}^{2}),\zeta_{v^{y}_{e,l}}\sim\mathcal{N}(0,\sigma_{v^{y}_{e,l}}^{2})$ and $\zeta_{{v^{z}_{e,l}}}\sim\mathcal{N}(0,\sigma_{v^{z}_{e,l}}^{2})$, respectively. Again, assuming mutually independent state prediction errors, we have $\boldsymbol{\zeta}_{l}\sim\mathcal{N}(0,\mathbf{Q}_{\mathbf{s}})$ with $\mathbf{Q}_{\mathrm{s}}=\mathrm{diag}\{\sigma_{x^{e}_{l}}^{2},\sigma_{y^{e}_{l}}^{2},\sigma_{z^{e}_{l}}^{2},\sigma_{v^{x}_{e,l}}^{2},\sigma_{v^{y}_{e,l}}^{2},\sigma_{v^{z}_{e,l}}^{2}\}$. Moreover, $\mathbf{F}\in\mathbb{R}^{6\times 6}$ is the state transition matrix. In the prediction model, we make the assumption of constant-velocity movement between adjacent CPIs and express the state transition matrix accordingly as
$\left.\mathbf{F}=\left[\begin{array}[]{cc}\mathbf{I}_{3}&\Delta t\mathbf{I}_{3}\\ \mathbf{0}_{3}&\mathbf{I}_{3}\end{array}\right.\right].$ (28)

### v2

@@ Abstract
Integrated sensing and communications (ISAC), secure communications, near field, unmanned aerial vehicle (UAV), 3D velocities sensing
@@ I Introduction
With the fast development of the low-altitude economy, unmanned aerial vehicles (UAVs) have received significant attention for potential usage in military, civilian, and commercial applications [1]. Particularly, benefiting from the high mobility and the low operation cost, UAVs are anticipated to play significant roles in future wireless networks [2]. In contrast to the terrestrial communications, the ground-to-air communications have higher probability to be dominated by the line-of-sight (LoS) links [3, 4], which is beneficial for improving the communication quality. Moreover, the UAV’s locations can be flexibly adjusted to adapt to the time-varying communications environment, which helps to construct preferable channel conditions in real time.
Despite the above advantages, one prominent challenge for facilitating UAV communications is the potential eavesdropping risk posed by the inherent broadcast nature of the LoS-dominated channels [5]. As such, it is critical to develop physical layer security (PLS) technologies to provide flexible security services for safeguarding UAV communications. Thanks to the capability of the multiple-input multiple-output (MIMO) technique for steering the signals toward the desired directions, the secrecy rate at the receivers can be enhanced via effective beamforming [6, 7]. However, since the eavesdroppers are usually non-cooperative nodes, it is impractical to obtain the wiretap channel state information (CSI) with the conventional pilot-based estimation approach [8]. In addition, due to the high mobility of the UAV, the CSI tracking is a tough task to guarantee stable communication performance while meeting the strict latency requirement. Recently, the concept of sensing-aided communication, which leverages the integrated sensing and communications (ISAC) techniques [9, 10], has emerged for obtaining the communication/wiretap CSI and thereby assisting the subsequent beamforming design [11, 12]. This approach eliminates the requirement for transmitting pilot signals, and instead estimating the CSI based on the distance and direction parameters obtained from echo signals reflected by legitimate users/eavesdroppers [13].
@@ I-B Motivations and Contributions
  • For communications, near-field propagation enables the unique beamfocusing capability, i.e., targeting the signal on a specific location in the near field instead of a specific direction in the far field. This provides enhanced flexibility to prevent legitimate information leakage and thus improve anti-eavesdropping performance, which is beneficial for secure UAV communications.
@@ I-C Organization and Notations
Notations: Vectors and matrices are denoted by boldface lowercase and boldface capital letters, respectively. $\mathbb{C}^{M\times N}$ denotes the space of all $M\times N$ matrices with complex entries. $\mathbf{A}^{*}$, $\mathbf{A}^{T}$ and $\mathbf{A}^{H}$ denote the conjugate, transpose, and conjugate transpose of the matrix $\mathbf{A}$, respectively. $\mathbf{A}_{i,j}$ returns the entry in the $i$-th row and $j$-th column of matrix $\mathbf{A}$. $\mathrm{tr}(\mathbf{A})$ and $\mathrm{rank}(\mathbf{A})$ denote the trace and the rank of matrix $\mathbf{A}$. $\odot$ is the Hadamard product. $\|\cdot\|$ denotes the norm of the vector. $\|\cdot\|_{F}$ denotes the Frobenius norm of a matrix. $\mathcal{CN}\left(u,v\right)$ denotes a Gaussian distribution with mean $m$ and variance $v$. $|\cdot|$ denotes the absolute value of a complex scalar. $\mathbf{I}_{M}$ denotes the $M\times M$ identity matrix. $\mathbf{0}_{M}$ denotes the $M\times M$ zero matrix. $x]^{+}$ stands for $\max\{x,0\}$.
@@ II System Model and Problem Formulation
As shown in Fig. 1, we consider a near-field [^1] We assume that both the C-UAV and E-UAV are in the near-field region during the whole flight. UAV communication system that consists of one ISAC GBS, one C-UAV receiving signals from the GBS, and one E-UAV wiretapping the GBS information. Thanks to few scatterers in the air-to-ground communication environment, we assume the LoS-dominant case for channels from the GBS to the C-UAV and the E-UAV [17, 22]. The GBS is deployed with a rectangular uniform planar array (UPA) that consists of $M=M_{x}\times M_{y}$ antennas with the spacing $d$, while the C-UAV and the E-UAV are both equipped with a single antenna. The index of the $m$-th GBS antenna is given by $m=aM_{y}+b$, where $a=0,\dots,M_{x}-1$ and $b=1,\dots,M_{y}$. We assume that the GBS operates in the full-duplex mode, which allows it to transmit and receive signals simultaneously. With the aid of advanced self-interference mitigation techniques [33], we assume that the full-duplex GBS can realize perfect self-interference mitigation. The bandwidth of the system, denoted by $B$, is related to the symbol duration $T_{s}=\frac{1}{B}$. Let $N$ denote the number of symbol durations in a single coherent processing interval (CPI) with the duration of $\Delta t$ for the near-field sensing, during which the UAVs’ location and velocity are assumed to be unchanged. We set a Cartesian coordinate system for which the origin is chosen as the center of the GBS UPA and the $x$ and $y$ axes are aligned with the column and row of the UPA, respectively, while the $z$ axis points up to complete a right handed coordinate frame. As such, the coordinate of the UPA center is given by $\mathbf{q}_{b}=[0,0,0]^{T}$, while that of the $m\mathrm{\text{-}th}$ antenna is denoted by $\mathbf{q}_{m}=[x_{m},y_{m},0]^{T}$. In the $l\in\{1,...,L\}\mathrm{\text{-}th}$ CPI, the position and the velocity of the E-UAV are denoted by $\mathbf{q}_{e,l}=\left[x^{e}_{l},y^{e}_{l},z^{e}_{l}\right]^{T}$ and $\mathbf{v}_{e,l}=[{v}^{x}_{e,l},{v}^{y}_{e,l},{v}^{z}_{e,l}]^{T}$, respectively, and the position of the C-UAV is represented by $\mathbf{q}_{c,l}=\left[x^{c}_{l},y^{c}_{l},z^{c}_{l}\right]^{T}$. Given that the E-UAV trajectory is unknown at the GBS, the GBS sends artificial noise (AN) signals to the E-UAV for both sensing and jamming purpose. Then, the transmitted signal $\mathbf{x}_{l}(n)$ of the GBS in the time slot $n$ of the $l$-th CPI is given by
@@ II-B Near-Field Communication Signal Model
where $r^{c}_{l},\theta^{c}_{l},\phi^{c}_{l}$ represent the distance, azimuthal angle, and elevation angle between the C-UAV and the GBS UPA center in the $l\text{-}\mathrm{th}$ CPI, respectively. We assume that the C-UAV CSI, i.e., $\mathbf{h}^{bc}_{l}$, is known at the GBS by employing the near-field channel estimation methods [35]. Accordingly, the achievable data rate at the C-UAV in the $l$-th CPI is given by
@@
$\mathbf{h}^{be}_{l}=\frac{\beta_{0}}{r_{l}^{e}}\boldsymbol{\alpha}(r^{e}_{l},\theta^{e}_{l},\phi^{e}_{l}).$ (11)
@@ III E-UAV 3D Trajectory Tracking Scheme
where $\mathbf{X}_{l}\left(\mathbf{f}^{D}_{l}\right)=\left[\mathbf{G}_{l}(1)\mathbf{x}_{l}(1),\dots,\mathbf{G}_{l}(N)\mathbf{x}_{l}(N)\right]$, $\mathbf{G}_{l}(n)=\mathbf{A}(r^{e}_{l},\theta^{e}_{l},\phi^{e}_{l})\odot\mathbf{D}_{l,n}(\mathbf{f}_{l}^{D})$, and $\mathbf{Z}_{l}^{b}=\left[\mathbf{z}_{l}(1),\mathbf{z}_{l}(2),\ldots,\mathbf{z}_{l}(N)\right]$. As shown in Fig. 2, to facilitate the 3D velocities derivation, we denote the E-UAV’s velocity in the spherical coordinate system that is centered at the GBS UPA center, as $\tilde{\mathbf{v}}_{e,l}=\left[v_{e,l}^{r},v_{e,l}^{\theta},v_{e,l}^{\phi}\right]$, where $v_{e,l}^{r}$, $v_{e,l}^{\theta}$, and $v_{e,l}^{\phi}$ represent the radial velocity, the azimuthal velocity, and the elevation velocity, respectively. Notably, compared to the 2D velocity sensing, the 3D velocity sensing involves more variables, and thus leads to higher complexity. As shown in Fig. 3, the mathematical relationship between $\mathbf{f}_{l}^{D}$ and $\tilde{\mathbf{v}}_{e,l}$ is expressed as
@@
In Eq. (15), $v^{r}_{e,l,m}$, $v^{\theta}_{e,l,m}$, and $v^{\phi}_{e,l,m}$ denote the projection of the E-UAV’s radial velocity, azimuthal velocity, and elevation velocity along the line connecting the E-UAV and the $m$-th GBS antenna, respectively. Thus, we have
@@
where $\alpha_{r,m}$, $\alpha_{\theta,m}$ and $\alpha_{\phi,m}$ represent the included angles between the line connecting the $m$-th GBS antenna and the E-UAV, and the E-UAV’s radial, azimuthal, and elevation velocities, respectively. These angles are illustrated in Fig. 3.
@@
Thanks to the various values of $f_{l,m}^{\text{D}},\forall m$, the ML method [32] can be utilized for the estimation of $\tilde{\mathbf{v}}_{e,l}$ based on the relationship between $\mathbf{f}_{l}^{D}$ and $\tilde{\mathbf{v}}_{e,l}$ as shown in Eq. (15). Specifically, the ML optimization problem is formulated as follows:
@@
$\hat{\beta}=\frac{\operatorname{tr}\left(\mathbf{Y}^{b}_{l}\mathbf{X}_{l}^{H}\left(\tilde{\mathbf{v}}_{e,l}\right)\right)}{\|\mathbf{X}_{l}\left(\tilde{\mathbf{v}}_{e,l}\right)\|_{F}^{2}}.$ (18)
@@
$\hat{\mathbf{v}}_{e,l}=\arg\max_{\tilde{\mathbf{v}}_{e,l}}g\left(\tilde{\mathbf{v}}_{e,l}\right)=\frac{\left|\mathrm{tr}\left(\mathbf{Y}^{b}_{l}\mathbf{X}_{l}^{H}\left(\tilde{\mathbf{v}}_{e,l}\right)\right)\right|^{2}}{\|\mathbf{X}_{l}\left(\tilde{\mathbf{v}}_{e,l}\right)\|_{F}^{2}}.$ (19)
@@
$\times\left[d_{l,n,1}\left(\tilde{\mathbf{v}}_{e,l}\right)\frac{\partial v_{e,l,1}}{\partial v^{i}_{e,l}},\ldots,d_{l,n,M}\left(\tilde{\mathbf{v}}_{e,l}\right)\frac{\partial v_{e,l,M}}{\partial v^{i}_{e,l}}\right]^{T},\vphantom{\int}$
@@
where $\tilde{\mathbf{v}}_{e,l}^{(t)}$ and $\eta^{(t)}$ denote the updated value of $\tilde{\mathbf{v}}_{e,l}$ and the step size in the $t$-th iteration. The details of the gradient-descent algorithm is shown in Algorithm 1. In Algorithm 1, the complexity for calculating $\partial g\left(\tilde{\mathbf{v}}_{e,l}\right)/\partial v^{i}_{e,l}$ is given by $\mathcal{O}\left(M^{2}N\right)$. Denote $I_{\text{grad}}$ as the number of iterations for the gradient descent convergence. Then, the overall computational complexity of Algorithm 1 is $\mathcal{O}\left(I_{\text{grad}}M^{2}N\right)$.
@@ III-B1 State Measurement Error Model
Assuming that the measurement errors for each state are uncorrelated, we can construct the measurement error covariance matrix as $\mathbf{P}_{l}=\text{diag}\{\sigma^{2}_{r^{e}_{l}},\sigma^{2}_{{\theta}^{e}_{l}},\sigma^{2}_{{\phi}^{e}_{l}},\sigma^{2}_{v^{r}_{e,l}},\sigma^{2}_{v^{\theta}_{e,l}},\sigma^{2}_{v^{\phi}_{e,l}}\}$. It has been demonstrated in [37] that the measurement errors are inversely proportional to the signal-to-noise ratio (SNR) of the echo signals. Therefore, we have $\sigma^{2}_{r^{e}_{l}}=\frac{c_{r}}{\mathrm{SNR}_{l}}$, $\sigma^{2}_{\theta^{e}_{l}}=\frac{c_{{\theta}}}{\mathrm{SNR}_{l}}$, $\sigma^{2}_{\phi^{e}_{l}}=\frac{c_{{\phi}}}{\mathrm{SNR}_{l}}$, $\sigma^{2}_{v_{r,l}}=\frac{c_{v_{r}}}{\mathrm{SNR}_{l}}$ , $\sigma^{2}_{v_{\theta}}=\frac{c_{v_{\theta}}}{\mathrm{SNR}_{l}}$ and $\sigma^{2}_{v_{\phi,l}}=\frac{c_{v_{\phi}}}{\mathrm{SNR}_{l}}$, where $c_{r}$, $c_{{\theta}}$, $c_{{\phi}}$, $c_{v_{r}}$$c_{v_{\theta}}$ and $c_{v_{\phi}}$ are the scaling factors [38], while $\mathrm{SNR}_{l}$ is given by
@@ III-B2 State Prediction Model
where $\boldsymbol{\zeta}_{l}=\left[\zeta_{x^{e}_{l}},\zeta_{y^{e}_{l}},\zeta_{z^{e}_{l}},\zeta_{v^{x}_{e,l}},\zeta_{v^{y}_{e,l}},\zeta_{v^{z}_{e,l}}\right]^{T}$, $\zeta_{x^{e}_{l}},\zeta_{y^{e}_{l}},\zeta_{z^{e}_{l}},\zeta_{v^{x}_{e,l}},\zeta_{v^{y}_{e,l}}$, and $\zeta_{v^{z}_{e,l}}$ denote the prediction errors for $x^{e}_{l},y^{e}_{l},z^{e}_{l},v^{x}_{e,l},v^{y}_{e,l}$, and $v^{z}_{e,l}$ respectively, with $\zeta_{x^{e}_{l}}\sim\mathcal{N}(0,\sigma_{x^{e}_{l}}^{2}),\zeta_{y^{e}_{l}}\sim\mathcal{N}(0,\sigma_{y^{e}_{l}}^{2}),\zeta_{z^{e}_{l}}\sim\mathcal{N}(0,\sigma_{z^{e}_{l}}^{2}),\zeta_{v^{x}_{e,l}}\sim\mathcal{N}(0,\sigma_{v^{x}_{e,l}}^{2}),\zeta_{v^{y}_{e,l}}\sim\mathcal{N}(0,\sigma_{v^{y}_{e,l}}^{2})$ and $\zeta_{{v^{z}_{e,l}}}\sim\mathcal{N}(0,\sigma_{v^{z}_{e,l}}^{2})$, respectively. Again, assuming mutually independent state prediction errors, we have $\boldsymbol{\zeta}_{l}\sim\mathcal{N}(0,\mathbf{Q}_{s})$ with $\mathbf{Q}_{s}=\mathrm{diag}\{\sigma_{x^{e}_{l}}^{2},\sigma_{y^{e}_{l}}^{2},\sigma_{z^{e}_{l}}^{2},\sigma_{v^{x}_{e,l}}^{2},\sigma_{v^{y}_{e,l}}^{2},\sigma_{v^{z}_{e,l}}^{2}\}$. Moreover, $\mathbf{F}\in\mathbb{R}^{6\times 6}$ is the state transition matrix. In the prediction model, we make the assumption of constant-velocity movement between adjacent CPIs and express the state transition matrix accordingly as
[…]